OpenStudy (anonymous):

Suppose the earth were smooth, an you could wrap a 25,000 mile long metal belt snugly around it. Now let's say you lengthen the band by 10 feet, loosening it just a little. WHat would be the largest thing that could slither under the new band? An amoeba, a worm, a snake, or an alligator? What did the smartest person in the world answer?

OpenStudy (anonymous):

alligator

OpenStudy (radar):

10/5280=.001894 mile expansion. Calculate radius (before expansion).....then calculate radius with the expansion, and the difference of the radii will be the amount that can be slipped under the wire. lol

OpenStudy (radar):

Do you really want to know?

OpenStudy (anonymous):

the radius will increase by 10 / pi - approximately 3 ft

OpenStudy (earthcitizen):

senseless

OpenStudy (radar):

Lets see radius before expansion is C=2piror 25,000=2Pir, r=25,000/2Pi=3978.873577 miles.

OpenStudy (anonymous):

my oppio=nion.... If the circumference of the earth were 25000 miles, then the radius would be: 25000 = 2π r r = 25000/(2π) = 3978.8736 miles (approx.) If the circumference is increased by 10 ft, or by 0.0018939... miles, Then the radius would become 25000.0018939/(2π) = 3978.8739 miles (approx.) That would mean the band would be raised above the surface by 0.0003 miles or 1.584 feet (approx.) You could fit an alligator under the band !

OpenStudy (anonymous):

my answer :\

OpenStudy (radar):

radius after expansion is 25,000.001894/2Pi=3978.873878 miles, radius increase by .000333 mile or 1.589808 feet. The answer is amoeba, worm, snake and the alligator provided it was less than 1.589808 feet in height.lol

OpenStudy (anonymous):

pi * (d + x) - pi * d = 10 pj * x = 10 x = 10 / pi x approx 3 so radius increase will be about 1.5 - i was wrong above but alligator shld still pass under

OpenStudy (radar):

Yes, all of the above lol

OpenStudy (anonymous):

not quite senseless - i believe!!

OpenStudy (radar):

radius after expansion is 25,000.001894/2Pi=3978.873878 miles, radius increase by .000333 mile or 1.589808 feet. The answer is amoeba, worm, snake and the alligator provided it was less than 1.589808 feet in height.lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

actual the increasein radius is independent of the original value

OpenStudy (earthcitizen):

i see, my bad

OpenStudy (radar):

But you have it in your equation the first line.....d+10 that d is the original diameter thus the original radius.

OpenStudy (radar):

That post was for @joeywhite

OpenStudy (radar):

@joeywhite, I like your method, no messy conversions and real long numbers etc.

OpenStudy (anonymous):

@radar what i mean is its independent of the value of the diameter

OpenStudy (radar):

Yes, you are right, and as I said much more straight forward.

OpenStudy (anonymous):

O.o

OpenStudy (anonymous):

- i just got the radius confused with the diameter in my first post

OpenStudy (radar):

It was a fun question and I learned how to do it simpler.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

well thanks for the help ^-^

OpenStudy (anonymous):

yw - it was a crafty question in a way - you were given the circumference of the earth although you didn't need it to solve the problem